# Hyperparameter Optimization in Machine Learning

### v2

@@ 3.1 Grid Search
Perhaps the first formal description of grid search (GS) dates back to the 1930’s, appearing in a foundational work by fisher1936design on experimental design, under the name of factorial design. GS consists in evaluating $f(\lambda)$ over a predefined grid of hyperparameter configurations $\upsilon\subset\Lambda$. For each entry of $\lambda\in\Lambda=\Lambda_{1}\times\dots\Lambda_{m}$, the user defines a list of values:
$\bar{\lambda}^{i}=\{\bar{\lambda}^{i}_{j}\}_{j=1}^{n_{i}},\quad\bar{\lambda}^{i}_{j}\in\Lambda_{i},\quad\text{for }\;i\in[m].$
The search grid is then given by all the combinations on the Cartesian product generated by these $m$ sets, namely $|\upsilon|=\prod_{i=1}^{m}n_{i}$. The learning algorithm can be executed sequentially or in parallel on the resulting set of points. The hyperparameter configuration that minimizes the objective function $f$ over $\upsilon$ is returned as $\hat{\lambda}$; that is
$\hat{\lambda}\in\operatorname*{arg\,min}_{\lambda\in\upsilon}f(\lambda).$ (3.1)
GS is a simple memoryless method that can be used to explore low-dimensional hyperparameter spaces. It requires very little implementation effort and parallelization is straightforward. The number of total trial generated by GS, however, grows linearly in the size of the trial sets $\bar{\lambda}^{i}$ and is exponential in $m$, making the method impractical for even moderate values of $m$. In addition, while Boolean and categorical hyperparameters naturally lend themselves to an evaluation on a predefined grid, unbounded integer and real-valued hyperparameters require manually setting bounds and discretization.
@@ 3.2 Random Search
Perhaps conceptually the simplest procedure for HPO, random search (RS) appears already considered in the 50’ as a method for global optimization [brooks1958discussion]. RS consists in repeatedly drawing hyperparameter configurations from a predefined distribution $\upsilon$ over $\Lambda$ and it is readily parallelized. The incumbent solution after $K$ trials is given by
$\hat{\lambda}_{K}=\operatorname*{arg\,min}_{(\lambda_{1},\dots\lambda_{k})\sim\upsilon^{K}}f(\lambda_{i}),$ (3.2)
where the trial points $\lambda_{i}$ are drawn i.i.d. from $\upsilon.$ While in GS the total number of trial points is predetermined, one may keep on executing RS indefinitely. Termination conditions might involve approximate convergence checks, like lack of improvements after a certain amount of trials, or, much more frequently in HPO, after a budget is exhausted (see Section 2.4). For instance, one may run RS for 12 hours.
Even if in principle $\upsilon$ could be any distribution, when referring to random search for hyperparameter optimization, one typically implies the use of simple distributions (uniform or log-uniform distributions) that factorize independently along the configuration components, i.e. $\upsilon=\prod_{j=1}^{m}\upsilon_{i}$. For instance, if $\lambda$ contains, in order, a categorical choice between $K$ different neural architectures, the learning rate for stochastic gradient descent and a global dropout factor for regularization then $\upsilon$ could be set as
$\upsilon=\mathcal{U}\{[k]\}\times\mathrm{Log}\,\mathcal{U}(10^{-5},10^{-1})\times\mathcal{U}(0,0.5).$
where $\mathcal{U}\{\cdot\}$ and $\mathcal{U}(\cdot)$ are the discrete and continuous uniform distributions and $\mathrm{Log}\,\mathcal{U}(\cdot)$ is the log-uniform (or reciprocal) distribution on the indicated intervals. Using a distribution rather than a grid allows for a more natural treatment of a larger class of hyperparameters: RS does not require discretization of continuous spaces and also in principle allows also for distributions with non-compact supports.
@@
3.2.1 On the efficiency of random search (compared to grid search).
@@
In an influential article, bergstra_random_2012 advocate for using RS as the “default” baseline method for HPO instead of manual or grid search. They argue that while RS retains the advantages of GS, such as reproducibility, conceptual simplicity, very low implementation overhead and ease of parallelization, it is also more efficient in higher dimensional spaces, especially when the response function has a low effective dimensionality. The authors informally define this condition as when a subset of the hyperparameters accounts for most of the variation of the response function. For instance, for a bi-dimensional hyperparameter $\lambda=(\lambda_{1},\lambda_{2})$ it could be that $f(\lambda)\approx g(\lambda_{1})$, for some $g:\Lambda_{1}\to\mathbb{R}$.
One way to reason about the success of RS is to assume that there is a sub-region $\Lambda^{*}\subset\Lambda$ that leads to a positive outcome of the experiment, i.e., where $f$ is sufficiently small. Then, we can assess the odds that a run of RS “lands” at least one configuration in the good region by comparing the volume of such region with that of the full search space under random sampling [brooks1958discussion]. We can model the probability of success of random search after $K$ draws as follows:
$\mathbb{P}_{(\lambda_{1},\dots\lambda_{K})\sim\upsilon^{K}}\left(\hat{\lambda}_{K}\in\Lambda^{*}\right)=1-\left(1-\mathbb{P}_{\lambda\sim\upsilon}(\lambda\in\Lambda_{*})\right)^{K},$ (3.3)
@@
where $\mathbb{P}_{\upsilon}$ denotes the probability of sampling from $\upsilon$, typically uniform or log-uniform. Eq. (3.3) says that the minimum number of trials needed to assure a positive outcome with a given probability, say $95\%$, depends only on the probability of hitting $\Lambda_{*}$ under $\upsilon$. This can be independent of the search space dimension if the response function has a low effective dimensionality. This argument may explain the success of RS as a baseline in seemingly complex settings such as neural architecture search [li2019random], where repeated experimentation on a few well-known benchmark datasets has led to the development of well-tuned search spaces for neural architectures.
@@
For a general search space, we should expect $\mathbb{P}_{\upsilon}(\lambda\in\Lambda_{*})$ to decay exponentially in some $m^{\prime}\leq m$. For concreteness, consider the case where $\Lambda=[0,1]^{m}$ is the $m$-dimensional unit hypercube and $\Lambda^{*}=[0,1/2]^{m^{\prime}}\times[0,1]^{m-m^{\prime}}$ is the “good region”. The $m^{\prime}$-dimensional 1/2-hypercube represents the important hyperparameters, while the $m-m^{\prime}$-dimensional unit hypercube represents the unimportant ones, as these coordinates do not matter in hitting $\Lambda^{*}$. If $\upsilon=\mathcal{U}(\Lambda)$ is the uniform distribution over $\Lambda$, then $\mathbb{P}_{\upsilon}(\lambda\in\Lambda_{*})=2^{-m^{\prime}}$. Consequently, the number of trials needed to obtain a positive outcome with at least $95\%$ probability is
$K\geq\log(0.05)/\log(1-2^{-m^{\prime}}).$
@@
For $m^{\prime}=3$ one already needs $K\geq 23$ draws to achieve a $95\%$ probability of success. This illustrates that the number of trials needed grows exponentially in $m^{\prime}$. At the same time, this also shows that if $m^{\prime}\ll m$ RS is much more efficient compared to GS.
@@ 3.2.2 Unlucky runs
One potential disadvantage of RS is that its randomized nature means that “unlucky runs” may end up with entire regions of $\Lambda$ being under-explored or completely neglected. As a consequence, a single execution of random search may not necessarily provide sufficient information to meaningfully modify $\upsilon$ for another iteration of random search, should it be needed.
@@
$\mathbb{P}_{(\lambda_{1},\dots,\lambda_{2^{m}})\sim\upsilon^{2^{m}}}(\forall\,j\in[2^{m}],\;\;\exists\,k\in[2^{m}]\,\text{ s.t. }\,\lambda_{k}\in I_{j})=\prod_{j=1}^{2^{m}-1}(1-j/2^{m}).$
For example, for $m=3$ such probability is already $3/32$,. This means that, in expectation, RS with $8$ draws covers all the hypercubes only once every ten runs.
@@ 3.3 Quasi-random search
One basic requirement for a stochastic generator of hyperparameter configuration is consistency: we would like that $\lim_{k\to\infty}\min_{i\in[k]}f(\lambda_{i})=f(\lambda^{*})$. RS with a uniform distribution is consistent, while GS is not. However, unlike GS, random search for any fixed budget may generate a set of configurations that misses large portions of the search space. Intuitively, good sequences should cover evenly the search space. Also their projections onto subspaces of $\Lambda$ should be well spread, in order to speed up convergence when optimizing functions with low effective dimensionality – a property that GS misses. Quasi-random search methods seek to combine the advantages of grid and random search, with configuration-drawing strategies that combine randomness with deterministic steps.
Formalizations of these desiderata have been developed in the context of numerical integration and global optimization via the concepts of low dispersion and low discrepancy sequences [see e.g. sobol1979systematic, niederreiter1988low, dick2010digital]. bousquet2017critical study the usage of such quasi-random sequences in HPO, finding that they consistently outperform random search (with a uniform distribution) on a series of deep learning tasks; see also [tuningplaybookgithub, Chapter 7] The resulting methods are conceptually very similar to both GS and RS, being highly parallelizable, except that the set of hyperparameters $\lambda_{1},\dots,\lambda_{k}$ are generated according to more complex procedures. We review two possible choices below, assuming that the search space is an $m$-dimensional unit hypercube for simplicity. Note that this is typically not at all an hard limitation: starting with samples on the unit interval, one can easily derive samples for a variety of distributions with appropriate reparameterizations [see, e.g. devroye2006nonuniform].
@@ 3.3.1 Latin hypercube sampling (LHS)
LHS [mckay1979comparison] directly combines grid and random search, in that it divides the search space into a grid, but samples configurations at random within regions of the grid.
LHS works as follows: choose a sequence length $n>1$, which determines the number of hyperparameters generated by a run of the sampling process. Divide each $[0,1]$ interval of the unit hypercube into $n>1$ parts, denoting by $I_{s}=[(s-1)/n,s/n]$ the resulting intervals. This gives rise to $n^{m}$ hypercubes of volume $n^{-m}$ each. Then generate $m$ random permutations of $[n]$, denoted by $\{\sigma_{i}\}_{i=1}^{m}$, $\sigma_{i}:[n]\to[n]$. For $k\in[n]$, sample hyperparameter configurations as follows:
@@ 3.3.2 Sobol’ sequences
Sobol’ sequences [sobol1979systematic] are a collection of separate one-dimensional sequences, one for each axis: when constructing a $m+m^{\prime}$ dimensional sequence one can reuse the $m$-dimensional one for the first coordinates.
Let $j\in[m]$ be an axis, and let $\mathbb{F}_{2}=\{0,1\}$ with integer operations modulo $2$. For each integer $k$ – representing the iterate – let $k=\sum_{l}a_{l}(k)2^{j}$ be the binary expansion of $k$, where $a_{l}(k)\in\mathbb{F}_{2}$ are binary coefficients (e.g., for $k=3$, these would be $a_{0}(3)=1$, $a_{1}(3)=1$ and the rest 00’s). Denote by $a(k)$ the resulting vector. Let $C_{j}$ be a specific binary (infinite) “direction matrix” for axis $j$. Then the $j$-th coordinate of the $k$-th iterate of a Sobol’ sequence can be written as
$[\lambda_{k}]_{j}=\sum_{l}\frac{[C_{j}\cdot a(k-1)]_{l}}{2^{l+1}}$
where the matrix-vector multiplication $C^{j}\cdot a(k)$ is in $\mathbb{F}_{2}$ (i.e., the additions are modulo 2). The direction matrices $C^{j}$ are constructed starting from primitive polynomials of $\mathbb{F}_{2}$ and direction numbers, which must be carefully chosen to ensure good coverage of the sequences. We refer the reader to [Lemieux:1338331, Chapter 5] for details.
Note that, given the directions matrices, Sobol’ sequences are deterministic. It is common practice to obtain stochastic variants of such sequences by adding random shifts, i.e. $\lambda^{\prime}_{k}=\lambda_{k}+u$ (modulo the unit hypercube) with $u\sim\mathcal{U}([0,1]$ and by “scrambling” the binary representation vector $a(k)$, that is, randomly permuting its entries [owen1995randomly].
The bottom row of Figure 3.1 plots 4, 9 and 16 points Sobol sequences with random shift and scrambling on the unit square.
@@ 3.3.3 Other schemes
bousquet2017critical experimented also with Halton and Hammersley sequences, finding the latter (with scrambling and random shifting) to be overall robust and well performing on the tested deep learning scenarios. Other quasi-random low-discrepancy sequences are presented in [Lemieux:1338331]. LHS, Sobol’, Halton and Hammerley sampling are implemented in the stats.qmc module of SciPy.
Figure 3.1: Non-adaptive sampling methods on the unit square for different sample sizes: 4 (left), 9 (center) and 16 (right) points. Each row represents a different schema. Grid and Latin hypercube sampling (LHS) may require re-sampling every time we increase the sample size. Instead, for random and Sobol we can keep on adding points, “reusing” previous samples. For each plot, we also show projections of the samples onto the two axes. We can see how grid samples cover much fewer points on the projections w.r.t. the other sampling methods. We can also observe the “unlucky run” phenomenon of (uniform) random sampling, where after nine draws the upper-right quadrant is still unvisited. LHS and Sobol (with random shifts and scrambling) provide more even coverage of the space and its unidimensional projections.
@@
The techniques described in the previous chapter do not make any attempt to model the response function $f(\lambda)$: in other words, evaluations of $f$ on previously selected configurations do not influence the search. In contrast, model-based (MB) techniques construct an explicit surrogate model for $f(\lambda)$ – or for closely related quantities – based on the information collected from previous evaluations. The surrogate model is a first central component of model-based HPO methods. We will denote such model by
$\hat{f}:\Lambda\times\mathcal{X}\to\mathcal{Y},$ (4.1)
where $\mathcal{X}$ may represent additional inputs (e.g., the performance of partial executions of $A$, a baseline or the value of the incumbent solution) and $\mathcal{Y}$ is an appropriate output space. For instance, $\mathcal{Y}$ could be (a distribution over) $\mathbb{R}^{+}$ if $\hat{f}$ models directly $f$, or the interval $[0,1]$ if $\hat{f}$ predicts the probability that a configuration improves over a given baseline.
The landscape of model-based methods is strongly dominated by techniques falling under the umbrella of Bayesian optimization (BO), which we cover in Section 4.2. In BO, the surrogate model is a probabilistic mapping that also captures the uncertainty of the estimation; implementations include Gaussian processes (Section 4.2.2) and tree Parzen estimators (Section 4.2.5). Regardless of the particular implementation, one critical feature of surrogate models is that they should be much cheaper to evaluate than $f$ itself. BO methods do not locally minimize the surrogate function to select the next configuration to evaluate, but rather mediate it through a so-called acquisition function (AF), a second central component of a BO method and defined as a mapping:
$\alpha:\Lambda\times\hat{\mathcal{F}}\times\mathcal{X}\to\mathbb{R}$ (4.2)
where $\hat{\mathcal{F}}$ is an appropriate space of surrogate models, and, again, $\mathcal{X}$ collects auxiliary inputs, that is used to select the next configuration to evaluate $f$ on. Typically, $\alpha$ encodes a trade-off between exploration and exploitation:
  • exploration seeks to evaluate $f$ on points with high uncertainty, with the primary aim to improve the surrogate model;
  • exploitation seeks to leverage knowledge gathered so far to improve on the incumbent solution.
We discuss acquisition functions in Section 4.2.3.
We briefly touch upon other approaches to model-based optimization that do not directly fall in the BO category in Section 4.3. We conclude the chapter with a practical example that contrasts BO to random search 4.4 and an overall discussion in Section 4.5. In particular, we discuss parallelizability of model-based methods, a challenging practical bottleneck of MB-HPO methods (Section 4.5.1).

### v3

@@ 3.1 Grid Search
Perhaps the first formal description of grid search (GS) dates back to the 1930s, appearing in a foundational work by fisher1935design on experimental design, under the name of factorial design. GS consists in evaluating $f(\lambda)$ over a predefined grid of hyperparameter configurations. For a generic hyperparameter $\lambda^{(i)}$ with domain $\Lambda^{(i)}$, we define the discretized domain $\bar{\Lambda}^{(i)}$ as a finite subset of $\Lambda^{(i)}$. The search grid is the Cartesian product
$\bar{\Lambda}=\bar{\Lambda}^{(1)}\times\bar{\Lambda}^{(2)}\times\cdots\times\bar{\Lambda}^{(m)}\subset\Lambda.$
The learning algorithm is then executed sequentially or in parallel on the resulting set of tuples and the returned solution is
$\hat{\lambda}\in\operatorname*{arg\,min}_{\lambda\in\bar{\Lambda}}f(\lambda).$ (3.1)
GS is a simple memoryless method that can be used to explore low-dimensional hyperparameter spaces. It requires very little implementation effort and parallelization is straightforward. Unfortunately, if $d$ is the cardinality of the largest discretized domain, then $|\bar{\Lambda}|=O(d^{m})$, making the method impractical for even moderate values of $m$. In addition, while Boolean and categorical hyperparameters naturally lend themselves to an evaluation on a predefined grid, unbounded integer and real-valued hyperparameters require manually setting bounds and discretization.
@@
Figure 3.1: Non-adaptive sampling methods on the unit square for different sample sizes: 4 (left), 9 (center) and 16 (right) points. Each row represents a different schema. Grid and Latin hypercube sampling (LHS) require re-sampling every time we increase the sample size. Instead, for random and Sobol’ we can keep on adding points, “reusing” previous samples. For each plot, we also show projections of the samples onto the two axes. We can see how grid samples cover much fewer points on the projections w.r.t. the other sampling methods. We can also observe the “unlucky run” phenomenon of (uniform) random sampling, where after nine draws the upper-right quadrant is still unvisited. LHS and Sobol’ (with random shifts and scrambling) provide more even coverage of the space and its unidimensional projections.
@@ 3.2 Random Search
Random search (RS) was already considered in the 1950s as a method for global optimization [brooks1958discussion]. It is perhaps the conceptually simplest procedure for HPO. RS consists in repeatedly drawing hyperparameter configurations from a predefined distribution $\upsilon$ over $\Lambda$, and it is readily parallelized. The solution after $T$ trials is given by
$\hat{\lambda}(T)=\operatorname*{arg\,min}_{\lambda\in\{\lambda_{1},\dots\lambda_{T}\}}f(\lambda),$ (3.2)
where the trial points $\lambda_{1},\dots,\lambda_{T}$ are drawn i.i.d. from $\upsilon\in\mathcal{P}(\Lambda)$. While in GS the total number of trial points is predetermined, one may keep on executing RS indefinitely. Termination conditions might involve approximate convergence checks, like lack of improvements after a certain amount of trials, or after a given budget is exhausted (see Section 2.4). For instance, one may run RS for 12 hours.
In principle, $\upsilon$ could be any distribution. However, when referring to random search for hyperparameter optimization, one typically implies the use of simple distributions (e.g., uniform or log-uniform distributions) that factorize independently along the configuration components, i.e., $\upsilon=\prod_{i=1}^{m}\upsilon^{(i)}$. For instance, if $\lambda^{(i)}$ is a categorical choice between $N$ different neural architectures, then $\upsilon^{(i)}$ could be a categorical uniform distribution. Similarly, if $\lambda^{(i)}$ is the learning rate for stochastic gradient descent, then $\upsilon^{(i)}$ could be the log-uniform (or reciprocal) distribution on the interval $(10^{-5},10^{-1})$.
Using a distribution rather than a grid allows for a more natural treatment of a larger class of hyperparameters: RS does not require discretization of continuous spaces and also allows for distributions with non-compact supports.
@@
3.2.1 On the efficiency of random search
@@
In an influential article, bergstra_random_2012 advocate for using RS as the “default” baseline method for HPO instead of manual or grid search. They argue that while RS retains the advantages of GS, such as reproducibility, conceptual simplicity, very low implementation overhead and ease of parallelization, it is also more efficient in higher dimensional spaces, especially when the response function has a low effective dimensionality. The authors informally define this condition as when a subset of the hyperparameters accounts for most of the variation of the response function. For instance, for a bi-dimensional hyperparameter $\lambda=(\lambda^{(1)},\lambda^{(2)})$ it could be that $f(\lambda)\approx g(\lambda^{(1)})$, for some $g\mathrel{\mathop{\ordinarycolon}}\Lambda^{(1)}\to\mathds{R}$.
One way to reason about the success of RS is to assume that there is a sub-region $\Lambda^{*}\subset\Lambda$ that leads to satisfactory performance (i.e., where $f$ is sufficiently small). Then, we can assess the probability that RS returns a solution in this sub-region [brooks1958discussion], assuming that hyperparameter assignments are drawn i.i.d. from $\upsilon$:
$\mathbb{P}\left[\hat{\lambda}(T)\in\Lambda^{*}\right]=1-\left(1-\mathbb{P}[\lambda\in\Lambda^{*}]\right)^{T}.$ (3.3)
@@
Eq. (3.3) says that the minimum number of trials needed to assure a satisfactory outcome with a given probability, say $95\%$, depends only on the probability of hitting $\Lambda^{*}$ under $\upsilon$. This can be independent of the search space dimension if the response function has a low effective dimensionality. This argument may explain the success of RS as a baseline in seemingly complex settings such as neural architecture search [li2019random], where repeated experimentation on a few well-known benchmark datasets has led to the development of well-tuned search spaces for neural architectures.
@@
For a general search space, we should expect $\mathbb{P}[\lambda\in\Lambda^{*}]$ to decay exponentially in some $m^{\prime}\leq m$. For concreteness, consider the case where $\Lambda=[0,1]^{m}$ is the $m$-dimensional unit hypercube and $\Lambda^{*}=[0,1/2]^{m^{\prime}}\times[0,1]^{m-m^{\prime}}$ is the “good region”. The $m^{\prime}$-dimensional 1/2-hypercube represents the important hyperparameters, while the $m-m^{\prime}$-dimensional unit hypercube represents the unimportant ones, as these coordinates do not matter in hitting $\Lambda^{*}$. If $\upsilon$ is the uniform distribution over $\Lambda$, then $\mathbb{P}[\lambda\in\Lambda^{*}]=2^{-m^{\prime}}$. Consequently, the number of trials needed to obtain a positive outcome with at least $95\%$ probability is
$T\geq\log(0.05)/\log(1-2^{-m^{\prime}}).$
@@
For $m^{\prime}=3$ one already needs $T\geq 23$ draws to achieve a $95\%$ probability of success. This illustrates that the number of trials needed grows exponentially in $m^{\prime}$. At the same time, this also shows that RS is much more efficient compared to GS if $m^{\prime}\ll m$.
@@ 3.2.2 Unlucky runs
One potential disadvantage of RS, due to its randomized nature, is that entire regions of $\Lambda$ could be under-explored or completely neglected. As a consequence, a single execution of random search may not necessarily provide sufficient information to meaningfully modify $\upsilon$ for another iteration of random search, should it be needed.
@@
$\prod_{j=1}^{2^{m}-1}(1-j/2^{m}).$
For example, for $m=3$ such probability is already $3/32$,. This means that, in expectation, RS with $2^{3}=8$ draws covers all the hypercubes only once every ten runs.
@@ 3.3 Quasi-Random Search
One basic requirement for a stochastic generator of hyperparameter configuration is consistency: we would like that $\lim_{T\to\infty}\min_{k\in[T]}f(\lambda_{k})=f(\lambda^{*})$. RS with a uniform distribution is consistent, while GS is not. However, unlike GS, random search for any fixed budget may generate a set of configurations that misses large portions of the search space, as we just discussed. Intuitively, good sequences should cover evenly the search space. Also, their projections onto subspaces of $\Lambda$ should be well spread, in order to speed up convergence when optimizing functions with low effective dimensionality — a property that GS misses. Quasi-random search methods seek to combine the advantages of grid and random search, with configuration-drawing strategies that combine randomness with deterministic steps.
Formalizations of these desiderata have been developed in the context of numerical integration and global optimization via the concepts of low dispersion and low discrepancy sequences [see e.g. sobol1979systematic, niederreiter1988low, dick2010digital]. bousquet2017critical study the usage of such quasi-random sequences in HPO, finding that they consistently outperform random search (with a uniform distribution) on a series of deep learning tasks; see also [tuningplaybookgithub, Chapter 7]. The resulting methods are conceptually very similar to both GS and RS, being highly parallelizable, except that the set of hyperparameters $\lambda_{1},\dots,\lambda_{T}$ are generated according to more complex procedures. We review two possible choices below, assuming that the search space is an $m$-dimensional unit hypercube for simplicity. Note that this is typically not at all a hard limitation: starting with samples on the unit interval, one can easily derive samples for a variety of distributions with appropriate reparameterizations [see, e.g. devroye2006nonuniform].
@@ 3.3.1 Latin hypercube sampling (LHS)
LHS [mckay1979comparison] is a stratified sampling scheme that directly combines grid and random search, in that it divides the search space into a grid, but samples configurations at random within regions of the grid.
LHS works as follows: choose a sequence length $n>1$, which determines the number of hyperparameter configurations generated by a run of the sampling process. Divide each $[0,1]$ interval of the unit hypercube into $n>1$ parts, denoting by $I_{s}=[(s-1)/n,s/n]$ the resulting intervals. This gives rise to $n^{m}$ hypercubes of volume $n^{-m}$ each. Then generate $m$ random permutations of $[n]$, denoted by $\{\sigma_{i}\}_{i=1}^{m}$, $\sigma_{i}\mathrel{\mathop{\ordinarycolon}}[n]\to[n]$. For $k\in[n]$, sample hyperparameter configurations as follows:
@@ 3.3.2 Sobol’ sequences
Sobol’ sequences [sobol1979systematic] are a collection of separate one-dimensional sequences, one for each axis: when constructing an $m+m^{\prime}$ dimensional sequence one can reuse the $m$-dimensional one for the first coordinates. Furthermore, unlike LHS, Sobol’ sequences are “extendible”, meaning that if $\{\lambda_{1},\dots,\lambda_{T}\}$ is a Sobol’ sequence, one can add new points $\lambda_{T+1},\dots,\lambda_{T+T^{\prime}}$} such that $\{\lambda_{1},\dots,\lambda_{T+T^{\prime}}\}$ remains a Sobol sequence. This is, in practice, a useful property shared with random search, as it allows one to avoid committing to a predefined budget $T$, offering flexibility to add more points as needed, without invalidating previous runs.
Let $j\in[m]$ be an axis, and let $\mathbb{F}_{2}=\{0,1\}$ with integer operations modulo $2$. For each integer $k$, representing the iterate, let $k=\sum_{l\geq 0}a_{l}(k)2^{l}$ be the binary expansion of $k$, where $a_{l}(k)\in\mathbb{F}_{2}$ are binary coefficients (e.g., for $k=3$, these would be $a_{0}(3)=1$, $a_{1}(3)=1$ and the rest 0’s). Denote by $a(k)$ the resulting vector. Let $C_{j}$ be a specific binary (infinite) “direction matrix” for axis $j$. Then the $j$-th coordinate of the $k$-th iterate of a Sobol’ sequence can be written as
$\lambda_{k}]_{j}=\sum_{l}\frac{[C_{j}\cdot a(k-1)]_{l}}{2^{l+1}}$
where the matrix-vector multiplication $C^{j}\cdot a(k)$ is in $\mathbb{F}_{2}$ (i.e., the additions are modulo 2). The direction matrices $C^{j}$ are constructed starting from primitive polynomials of $\mathbb{F}_{2}$ and direction numbers, which must be carefully chosen to ensure good coverage of the sequences. We refer the reader to Lemieux:1338331 for details.
Note that, given the directions matrices, Sobol’ sequences are deterministic. It is common practice to obtain stochastic variants of such sequences. This is achieved by first adding random shifts, i.e. replacing $\lambda_{k}$ with $\lambda_{k}+u$ (modulo the unit hypercube), where $u$ is drawn uniformly from the interval $[0,1]$, and by then randomly permuting the entries of the binary representation vector $a(k)$ [owen1995randomly]. The bottom row of Figure 3.1 shows 4, 9, and 16 points randomized Sobol’ sequences on the unit square.
@@ 3.3.3 Other schemes
bousquet2017critical experimented also with Halton and Hammersley sequences and found that the latter (with random permutations and shifts) performed well on the tested deep learning scenarios. Like Sobol’ sequences, both Halton and Hammersley sequences can be extended. Lemieux:1338331 presents other quasi-random low-discrepancy sequences. LHS, Sobol’, Halton and Hammerley sampling are implemented in the stats.qmc module of SciPy.
@@
The techniques described in the previous chapter do not attempt to model the response function and are memoryless, i.e., the outcomes of previous evaluations do not influence the search. In contrast, model-based techniques use data gathered from past evaluations to build an explicit approximation of the objective function, known as a surrogate model, which is the central component of the family of HPO methods described in this chapter.
In general, the surrogate model takes as input a candidate configuration and, optionally, additional contextual information, such as the performance of partially trained models or the performance of the current best configuration (the incumbent). It then produces a prediction about how well the configuration is expected to perform. For example, the surrogate model might predict a distribution over performance scores. Details on the functional form of the surrogate depend on the specific instantiation of the model-based HPO method and will be discussed in the next sections. Regardless of the exact implementation, a key requirement of surrogate models is that they must be much cheaper to evaluate than the original objective function.
The landscape of model-based (MB) methods is strongly dominated by techniques falling under the umbrella of Bayesian optimization (BO), which we cover in Section 4.2. In BO, the surrogate model is a probabilistic mapping that also captures the uncertainty of the estimation of the response function. Common implementations include Gaussian processes (Section 4.2.2) and tree Parzen estimators (Section 4.2.5). BO methods do not select the next configuration by directly minimizing the surrogate model. Rather, they use a second mapping, called acquisition function, which mediates the selection process. Typically, the acquisition function aims to balance the selection of the next configuration between two competing objectives:
  • Exploration, which prioritizes evaluating the response function in regions where the surrogate model is highly uncertain, with the goal of gathering information to improve the surrogate model itself;
  • Exploitation, which focuses on selecting configurations that the surrogate model predicts to perform well, aiming to improve the current best solution.
We examine various acquisition functions and their roles in Section 4.2.3, and present a practical example that contrasts BO to random search in Section 4.2.6. In Section 4.3, we briefly touch upon other approaches to model-based optimization that do not directly fall under the BO category. We conclude the chapter with an overall discussion of model-based methods in Section 4.4. In particular, we discuss the parallelizability of model-based methods — a significant practical bottleneck of MB-HPO methods — in Section 4.4.1.